OpenStudy (anonymous):

uhmmm ya i dk how to do these still d(-2)= -3/2(-2)+5

OpenStudy (abb0t):

I think you already figured it out. It looks like you plugged in (-2) for whatever variable you are given in d(x) or d(t) etc.,//

OpenStudy (anonymous):

ya uhm i need the answer i know i plugged in the -2 but thats not what i was looking for

OpenStudy (abb0t):

Multiply and add.

jimthompson5910 (jim_thompson5910):

d(-2)= -3/2(-2)+5 d(-2)= (-3/2)(-2/1)+5 ... rewrite -2 as -2/1 d(-2)= (-3*(-2))/(2*1)+5 ... multiply straight across d(-2)= 6/2+5 ... multiply I'll let you finish. Remember to follow PEMDAS

OpenStudy (anonymous):

17?

jimthompson5910 (jim_thompson5910):

how are you getting that?

OpenStudy (abb0t):

How did you get 17?

OpenStudy (anonymous):

idkkkkkk omggggod

OpenStudy (abb0t):

\[\frac{-3}{2}(-2)+5\]

OpenStudy (abb0t):

The (-2) and (2) cancel out to give you (-1). (-3)(-1) + 5 = ??